Suppose $1 = 10.5 pesos in New York and $1 = 9.6 Pesos in Mexico City. If you had $10,000 using arbitrage, your profits would be?
Answer #1

$10,000 = 9.6*10,000 Pesos or 96,000 Pesos in Mexico

$10,000 = 10.5*10,000 Pesos or 105,000 Pesos in New York

In Mexico 105000 Pesos = 105000/9.6 = $10937.5

In an act of arbitrage, if the Pesos are sold in Mexico it will yield $10,937.5

Therefore Profits from arbitrage = $10,937.5 - $10,000 = $937.50